Please help me solve:
If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take for both of them to paint the house together?



Answer by ewatrrr(24785)


Hi,
Sally can paint a house in 4 hrs, and John can paint the same house in 6 hrs
Let x represent the time it takes them together. Per hour the Equalizer
1/4 + 1/ 6 = 1/x Mulitplying each term by 12x so as all denominators = 1
3x + 2x = 12
5x = 12
x = 12/5 Or 2.4 hrs, the time it takes them together (2hr & 24min)
Note: the time painting the house together is always considerably less than
the time it takes one of them to do it alone. In this case Sally can paint the
house ALONE in 4 hrs..an answer of 3hrs and 12 min Together...NO
that would mean John would take 16hr to paint the house alone ..
1/4 + 1/16 = 1/x , 4x + x = 16, 5x = 16, x = 3.2 Or 3hrs and 12 min.